Total cost assigned to products Xtreme Pathfinder Total % of % of Amount Total Amount Amount Total Amount Amount Activity-Bas3. Prepare a quantitative comparison of the traditional and activity-based cost assignments. Complete this question by enteriSmoky Mountain Corporation makes two types of hiking boots-the Xtreme and the Pathfinder. Data concerning these two product l3. Prepare a quantitative comparison of the traditional and activity-based cost assignments. Complete this question by enteri3. Prepare a quantitative comparison of the traditional and activity-based cost assignments. Complete this question by enteri
Answer #1

For Smoky Mountain Corporation

(1) Product Margin ( from (2) and (3) below)

Particulars Xtreme Pathfinder Total
Traditional Cost System (in %) 7.6 11.8 10.5
Activity based Cost System (in %) 4.3 13.3 10.5

(2) Traditional Cost System

Particulars Particulars % Xtreme Particulars % Pathfinder Total
1.Total sales 23000 u@$124p.u 31.6 $2852000 70000u@$88p.u 68.4 $6160000 $9012000
2.Variable Cost
Direct Materials 23000u@$63.4p.u 28.2 $1458200 70000u@$53p.u 71.8 $3710000 $5168200
Direct Labor* 23000u@$14.4p.u 37.2 $331200 70000u@$8p.u 62.8 $560000 $891200
Variable Overheads** 23000u*1.8hrs*$10.92 37.2 $452324 70000u*1hr*$10.92 62.8 $764636 $1216960
Total (2) $2241724 $5034636 $7276360
3.Contribution(1-2) 21.4 $610276 18.3 $1125364 $1735640
4.Fixed Costs(from WN-2)*** $394120 $394120 $788240
5.Profit 7.6 $216156 11.8 $731244 $947400

*Total labor hours = Hours for Xtreme+ Hours for Pathfinder=1.8hrs*23000u+1hr*70000u=111400 hrs

**Total Manufacturing overheads =$2005200

Total Variable manufacturing overheads =$2005200-$788240=$1216960

Variable manufacturing overheads per hour = $1216960/111400 hrs=$10.92p.hr

Also rounding off difference is distributed equally

***Other Costs i.e., Fixed Costs =$48240+$740000=$788240

(3) Activity based Cost System

Particulars Particulars % Xtreme Particulars % Pathfinder Total
1.Total sales 23000 u@$124p.u 31.6 $2852000 70000u@$88p.u 68.4 $6160000 $9012000
2.Variable Cost
Direct Materials 23000u@$63.4p.u 28.2 $1458200 70000u@$53p.u 71.8 $3710000 $5168200
Direct Labor* 23000u@$14.4p.u 37.2 $331200 70000u@$8p.u 62.8 $560000 $891200
Variable Overheads From WN-1(3) 44.4 $540960 From WN-1(3) 55.6 $676000 $1216960
Total (2) $2330360 $4946000 $7276360
3.Contribution(1-2) 18.3 $521640 19.7 $1214000 $1735640
4.Fixed Costs $394120 $394120 $788240
5.Profit 4.3 $127520 13.3 $819880 $947400

WN-1

Particulars Amount ($) Xtreme Pathfinder
1.Supporting Direct Labor[$712960/111400hrs] 6.4 p.hr $6.4*41400hrs=$264960 $6.4*70000hrs=$448000
2.Batch Setups [$504000/420 set ups] 1200 p. set up $1200*230=$276000 $1200*190=$228000
3.Variable Manufacturing Overheads[1+2] $540960 $676000
4.Sustaining Costs i.e., Fixed Costs $370000 $370000
5.Other Costs*(WN-2) $24120 $24120
6.FixedCosts [4+5] $394120 $394120

WN-2

1.Sustaining costs are in the nature of fixed costs, also in the absence of information it is assumed that other costs are also apportioned between the two products equally.